OpenStudy (anonymous):

The Hudson River flows at a rate of 3mph. A patrol boat travels 60 mi upriver and returns (another 60mi ) What is the speed of the boat in still water? can somebody there help me to do this?

OpenStudy (anonymous):

wait, how long does the boat take total trip?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are missing info. We need the total time of the trip

OpenStudy (anonymous):

The Hudson River flows at a rate of 3mph. A patrol boat travels 60 mi upriver and returns (another 60mi ) in a total time of 9hr.What is the speed of the boat in still water? can somebody there help me to do this?

OpenStudy (anonymous):

Alright, now we have the time. Let S = speed of patrol boat in still water and T = time traveled going up river . (S-3)T = 60 (S+3)(9-T) = 60 . (S+3)(9-T) = 60 9S+27-ST-3T = 60 . (S-3)T = 60 T = 60/(S-3) . 9S+27-ST-3T = 60 9S+27-S(60/(S-3))-3(60/(S-3)) = 60 9S(S-3)+27(S-3)-S(60)-3(60) = 60(S-3) 9S^2-27S+27S-81-60S-180 = 60S-180 9S^2-27S+27S-81-60S = 60S 9S^2-81-60S = 60S 9S^2-81-120S = 0 9S^2 - 120S - 81 = 0 3S^2 - 40S - 27 = 0 . you must either factor or apply the quadratic formula. In this case, you can only apply the quadratic formula. It will give you two answers -- one negative and one positive. Since, a negative speed won't make sense here, the only one that makes sense is the positive answer. S = 13.98 mph .

OpenStudy (anonymous):

thank you